# Study of quantum phase transition and entanglement in coupled top systems with standard and nonstandard symmetries under Floquet formalism

### v1

@@ IV Classical dynamics of the effective Hamiltonian of the CKT
Following the above prescription, we present a rigorous derivation of the equation of motion of only one component of the angular momentum, say $J_{x1}$. Others are derived identically, and we just present the final form of them. Equation of motion of angular momentum operator ${J}_{x1}$ can be obtained as:
$\frac{d{J}_{x1}}{dt}$ $=-\imath[{J}_{x1},{H}_{\rm eff}]$
$=-\frac{\imath}{2j}\biggl{(}\kappa_{1}\,[{J}_{x1},{J}_{z1}^{2}]+2\varepsilon\,[{J}_{x1},{J}_{z1}]{J}_{z2}\biggr{)}$
$=-\frac{\imath}{2j}\biggl{\{}\kappa_{1}\,\bigl{(}[{J}_{x1},{J}_{z1}]{J}_{z1}+{J}_{z1}[{J}_{x1},{J}_{z1}]\bigr{)}$
@@
Now, divide the above equation with $j$ and obtain the form:
@@
The classical Hamiltonian $H^{\rm cl}_{\rm eff}$ corresponding to the above equation of motion is obtained by dividing the quantum effective Hamiltonian $H_{\rm eff}$ by the spin $j$ and set the limit $j\rightarrow\infty$. Thus we find $H_{\rm eff}^{\rm cl}$ in terms of the rescaled angular momenta as:
$H^{\rm cl}_{\rm eff}$ $=\lim_{j\rightarrow\infty}\frac{H_{\rm eff}}{j}=\left(\Omega_{1}\,X_{1}+\frac{1}{2}\kappa_{1}\,Z_{1}^{2}\right)$
$+$ $\left(\Omega_{2}\,X_{2}+\frac{1}{2}\kappa_{2}\,Z_{2}^{2}\right)+\varepsilon\,Z_{1}Z_{2}.$ (10)
We exploit the constraints $X_{i}^{2}+Y_{i}^{2}+Z_{i}^{2}=1$, where $i=1,2$, and can reduce the degrees of freedom from six to four. For this dimension reduction, we parameterize the angular momentum variables as $X_{i}=\sin\theta_{i}\,\cos\phi_{i},\,Y_{i}=\sin\theta_{i}\,\sin\phi_{i},$ and $Z_{i}=\cos\theta_{i}$, where $Z_{i}$ and $\phi_{i}$ become canonically conjugate variables for the $i$-th top. In terms of this new parameterization, the classical Hamiltonian becomes
$H_{\rm eff}^{\rm cl}=$ $\Omega_{1}\,\sqrt{1-Z_{1}^{2}}\,\cos\phi_{1}+\Omega_{2}\,\sqrt{1-Z_{2}^{2}}\,\cos\phi_{2}$
and the corresponding Hamilton’s equations of motion are derived as:
We have studied the effective Hamiltonian of the CKT system at two limits: (1) torsion-free limit $\kappa_{1}=\kappa_{2}=0$, this is the well-known Feingold-Peres (FP) model [42]; and (2) nonzero torsion case, i.e., $\kappa_{i}\neq 0$ with $i=1,2$. The latter case can further be classified as $\kappa_{1}=\kappa_{2}$ and $\kappa_{1}\neq\kappa_{2}$. For the unequal torsions, we only consider the case $\kappa_{1}=-\kappa_{2}$, and the quantum spectrum corresponding to this system follows one of the nonstandard symmetries.
IV.1 FP model (torsion-free case)
Figure 1: The dynamics of the coupled top are projected on the phase space of the first top at the FP limit. The FP limit is obtained by setting $\kappa_{1}=\kappa_{2}=0$. We set $\Omega_{1}=\Omega_{2}=1.0$. Left panel: Coupling $\varepsilon=0.8$; and Right panel: $\varepsilon=1.3$.
Figure 1 presents the phase space dynamics of the coupled top at the FP limit. Here, we fix $\Omega_{1}=\Omega_{2}=1.0$ and consider two different coupling strengths $\varepsilon=0.8$ (Left panel) and $\varepsilon=1.3$ (Right panel). We have chosen these two coupling strengths because, at $\varepsilon=1.0$, the phase space dynamics of the FP system show a transition, which is clear from the appearance of substructures at the center.
IV.2 Nonzero torsions case
We now present the classical dynamics of the effective Hamiltonian $H^{\rm cl}_{\rm eff}$ of the couple top. Here again, we set the parameters $\Omega_{1}=\Omega_{2}=1.0$ and consider the same two coupling cases, $\varepsilon=0.8$ and $1.3$, to show the effect of the nonlinear torsional terms. In Fig. 2, we consider two cases of torsions: at the top panels, two torsional terms are equal in magnitude and sign, i.e., $\kappa_{1}=\kappa_{2}$; and at the bottom panels, we consider torsional terms with opposite signs, but of the same magnitude, i.e., $\kappa_{1}=-\kappa_{2}$. For the $\kappa_{1}=\kappa_{2}$ case, presented in the top panels, the phase space dynamics are qualitatively similar for both the coupling strengths. This behavior is markedly different from the FP case, as discussed above. However, for the latter case, when $\kappa_{1}=-\kappa_{2}$, the dynamics are very much different from the case with the same value of the torsions. Here, we see that, at the center part of the phase space of the first top, the trajectories are diffusing from one stable island to another. Here, we have not plotted the Poincare section; instead, we have just projected the phase space trajectories of the coupled top residing on the four-dimensional phase space to the phase space of the first top having dimension two. Therefore, this diffusion of trajectories from one island to another happens through higher dimensional phase space. We also see the same behavior in other parts of the phase space.
V Symmetry properties of the effective Hamiltonian of the CKT
V.1 FP model (torsion-free case)
For the FP model, the effective Hamiltonian of the CKT becomes
$H_{\rm eff}=H_{\rm FP}=\Omega_{1}J_{x1}+\Omega_{2}J_{x2}+\frac{\varepsilon}{j}J_{z1}J_{z2}$ (13)
$U_{0}=e^{-\imath\pi J_{x1}}\otimes e^{-\imath\pi J_{x2}},$ (14)
$C=e^{\imath\alpha}e^{-\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}},$ (15)
which gives $C\,H_{\rm FP}\,C^{\dagger}=-H_{\rm FP}$ or $\left(C\,H_{\rm FP}+H_{\rm FP}\,C\right)=0$. This implies that ${\rm Tr}\,(H_{\rm FP})=0$. The phase factor $e^{i\alpha}$ is not playing any role here to show this symmetry of $H_{\rm FP}$ under $C$ because its complex conjugate factor from $C^{\dagger}$ part will trivially cancel it. However, this innocent-looking phase factor will become important while classifying the system under different symmetry classes. The operator, $C$, is called the chirality operator. In the case of the integer spin $C^{2}=\mathbbm{1}$. This property of $H_{\rm FP}$ implies the presence of nonstandard symmetries different from the standard Wigner-Dyson threefold symmetry classes. In addition, the FP Hamiltonian has a nonunitary symmetry like the time-reversal symmetry. We denote the time-reversal operator $\mathcal{T}$, which, on a standard basis, flips the sign of the $J_{y}$ operator but keeps the other two angular momenta invariant. Therefore, $H_{\rm FP}$ also satisfies $\mathcal{T}\,H_{\rm FP}\mathcal{T}^{-1}=H_{\rm FP}$. Depending on whether the chirality operator $C$ commutes with $\mathcal{T}$, we get different classes of symmetries for $H_{\rm FP}$. If the spin $j$ is integer, then we set $\alpha=0$ and find that
$\mathcal{T}\,C\,\mathcal{T}^{-1}$ $=\mathcal{T}\left[e^{-\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}}\right]\mathcal{T}^{-1}$
$=e^{\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}}$
$=e^{2\imath\pi J_{z1}}\,e^{-\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}}$
$=C$ (16)
where we use $\mathcal{T}\,\imath\,\mathcal{T}^{-1}=-\imath$; and for integer $j$, the term $e^{2\imath\pi J_{z1}}$ becomes identity matrix. Thus we find $\mathcal{T}\,C\,\mathcal{T}^{-1}=C$. However, for the half-integer spin $e^{2\imath\pi J_{z1}}\neq\mathbbm{1}$ and hence $\mathcal{T}\,C\,\mathcal{T}^{-1}\neq C$. For this case, the phase $\alpha$ can be tuned to get a $C$ which commutes with $\mathcal{T}$. For any arbitrary phase $\alpha$, we have,
$\mathcal{T}\,C\,\mathcal{T}^{-1}$ $=\mathcal{T}\left[e^{\imath\alpha}e^{-\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}}\right]\mathcal{T}^{-1}$
$=e^{-\imath\alpha}e^{\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}}$
$=e^{-\imath 2\alpha}\,e^{\imath 2\pi J_{z1}}\,\left[e^{\imath\alpha}\,e^{-\imath\pi J_{z1}}\,\otimes e^{-\imath\pi J_{y2}}\right]$
$=e^{\imath 2(\pi J_{z1}-\alpha)}\,C.$ (17)
$C=e^{\imath\pi j}e^{-\imath\pi J_{z1}}\otimes e^{-\imath\pi J_{y2}}$ (18)
V.2 Nonzero torsions case
For the nonzero torsions, first consider $\kappa_{1}=\kappa_{2}=\kappa$ and $\Omega_{1}=\Omega_{2}=\Omega$. Then the effective Hamiltonian $H_{\rm CT}$ of the coupled top becomes:
$=H_{\rm FP}+\frac{\kappa}{2j}\left(J_{z1}^{2}+J_{z2}^{2}\right)$
$\equiv H_{\rm FP}+H_{\rm NL},$ (19)
$=\frac{1}{j}\sum_{m=1}^{j}m^{2}\left(\kappa_{1}+\kappa_{2}\right)$
$=\frac{1}{6}(j+1)(2j+1)\left(\kappa_{1}+\kappa_{2}\right).$ (20)
The above relation shows that the trace is zero when $\kappa_{1}=-\kappa_{2}=\kappa$. Therefore, the CT Hamiltonian with trace zero is of the form
We have observed very different classical dynamics for this particular case of torsion with opposite signs. However, the Hamiltonian $H_{\rm CT}$ still does not have the chiral symmetry under the transformation of $C$. This result implies that there exists a different chirality operator, say $C^{\prime}$, under which $C^{\prime}\,H_{\rm CT}\,C^{\prime\dagger}=-H_{\rm CT}$. One point is to be noted that now the Hamiltonian $H_{\rm CT}$ is not symmetric under permutation. We exploit this fact and obtain $C^{\prime}=PC$, where $P$ is the permutation operator. One can now easily check that $C^{\prime}\,H_{\rm CT}\,C^{\prime\dagger}=P\,C\,H_{\rm CT}\,C^{\dagger}P^{\dagger}=-H_{\rm CT}$ or $\left(C^{\prime}\,H_{\rm CT}+H_{\rm CT}\,C^{\prime}\right)=0$. Since the chirality operator $C$ is time-reversal symmetric, one can show that the chirality operator $C^{\prime}$ for the couple top Hamiltonian is also time-reversal symmetric. Moreover, for $C^{2}=\pm\mathbbm{1}$, the other chirality operator also satisfies the same property, i.e., $C^{\prime 2}=\pm\mathbbm{1}$. Therefore, the CT Hamiltonian can also be classified into two same chiral symmetry classes, BDI and CI [32, 31]. Besides, for the nonzero trace case, one can not find a chirality operator under which the CT Hamiltonian will show the symmetry property. Therefore, this Hamiltonian belongs to the standard symmetry class.
VI Entanglement Calculation
We now study the entanglement property of the coupled-top system represented by the Hamiltonian $H_{\rm CT}$ for zero (FP case) and nonzero torsion cases. For the latter case, we consider $\kappa_{1}=\kappa_{2}$ and $\kappa_{1}=-\kappa_{2}$. We study entanglement in the system by calculating the entanglement of the ground and the most excited states of the Hamiltonian $H_{\rm CT}$. Since these are pure states, we calculate von Neumann entropy $S_{V}$ of the reduced density matrix (RDM) corresponding to the individual top as a measure of entanglement. The von Neumann entropy is defined as follows:
$S_{V}$ $=-{\rm Tr}_{1}[\rho_{1}\,\ln\rho_{1}]$
$=-{\rm Tr}_{2}[\rho_{2}\,\ln\rho_{2}],$ (22)
where $\rho_{1}$ and $\rho_{2}$ are the RDMs corresponding to the individual tops. In the eigenbasis of the RDMs, the von Neumann entropy can be expressed as:
$S_{V}=-\sum_{i}\lambda_{i}\ln\lambda_{i},$ (23)
where $\lambda_{i}$’s are the eigenvalues of the RDMs.
     Ground State            Excited State
Figure 3: Showing the variation of entanglement (von Neumann entropy) with coupling strength. The left column presents the results for the ground state, and the right column for the most excited state.
In Fig. 3, we present the results of entanglement calculation. Here, the variation of the von Neumann entropy is shown as a function of the coupling strength $\varepsilon$ for $j=10$ for the ground state and the most excited state. Figure 3(a) shows the result for the ground state energy of the FP model (i.e., $\kappa_{1}=\kappa_{2}=0$). We find that, initially, $S_{V}$ increases linearly with $\varepsilon$ and attains a maximum value $S_{V}^{\rm max}\simeq 0.91$ at $\varepsilon=1.2$, and it then immediately starts falling with the increment of the coupling strength. Later, we show that this transition in entanglement is consistent with the quantum phase transition in the system. The entanglement falls till the coupling strength $\varepsilon\simeq 3.0$ and saturates at a value close to $\ln 2$ with some fluctuations. This saturation of $S_{V}$ around $\ln 2$ indicates permutation symmetry in the system, i.e., if we interchange the two tops, the Hamiltonian $H_{\rm CT}$ remains invariant. If we consider two non-identical tops by considering $\Omega_{1}\neq\Omega_{2}$, then the permutation symmetry will be broken, and consequently, the entanglement between the tops approaches zero for higher coupling strengths. In Fig. 3(b), we present the results for the most excited state. Here also, the transition in the entanglement happens at the same coupling strength $\varepsilon=1.2$. Interestingly, we do not see the effect of permutation symmetry on the excited state, and it becomes completely unentangled with $S_{V}=0$ for stronger coupling strengths.
In Figs. 3(c) and (d), the von Neumann entropy $S_{V}$ is presented for the non-zero torsion case with $\kappa_{1}=\kappa_{2}=1.0$. In this case, like the previous one, the entanglement of the ground state initially increases linearly with the coupling strength $\varepsilon$ and reaches a maximum $S_{V}^{\rm max}\simeq 1.0$ at $\varepsilon=2.0$. The entanglement then decreases and finally saturates at $\ln 2$ with some fluctuations for stronger coupling strengths $\varepsilon\gtrsim 4.0$. This saturation of entanglement at $S_{V}\simeq\ln 2$, also observed in the FP model, can be understood in the following way. At the strong coupling limit, the individual dynamics of the tops become almost negligible. The spin-spin coupling term in the Hamiltonian becomes the dominant term, and then the energy eigenstates approximately become $|m_{1}\rangle\otimes|m_{2}\rangle$, where $|m_{i}\rangle$’s are the eigenstates of $J_{z_{i}}$ and $m_{i}=-j,\dots,+j$. These eigenstates are unentangled states. However, due to the presence of permutation symmetry in the system, $|\psi_{m_{1},m_{2}}\rangle=\frac{1}{\sqrt{2}}\left(|m_{1}\rangle\otimes|m_{2}\rangle\pm|-m_{1}\rangle\otimes|-m_{2}\rangle\right)$ are also valid eigenstates with the same energy. The von Neumann entropy of these eigenstates $|\psi_{m_{1},m_{2}}\rangle$ equals to $\ln 2$. However, this nonzero entanglement of these states is not useful because this cannot be used as a resource for any quantum protocols. This is just an artifact of the underlying permutation symmetry in the system.
Figures 3(e) and (f) show the results for the non-zero torsion case with $\kappa_{1}=-\kappa_{2}=1.0$. This version of the couple top system is interesting because it follows one of the nonstandard symmetry classes [32, 31]. Since we consider integer spin here, the system follows the BDI symmetry class. In this case, the ground state entanglement also increases linearly and reaches its maximum $S_{V}^{\rm max}\simeq 0.8$ at $\varepsilon\simeq 1.0$ and suddenly starts decreasing. Here, the entanglement transition point is the same as the FP model. Like the previous two cases, the entanglement immediately goes down to $\sim\ln 2$ immediately after reaching the maximum, and then for $\varepsilon\simeq 3.0$, the system becomes unentangled. Interestingly, for $\varepsilon\gtrsim 3.0$, the von Neumann entropy increases slowly in a linear fashion with some fluctuations. The behavior of the excited state entanglement is almost similar to the ground state, except we observe that the excited state becomes unentangled or product state at $\varepsilon\simeq 3.0$ and remains as its for larger values of coupling strength.
Overall, we find that the entanglement transition in the ground state of the FP model and the nonzero torsional model with $\kappa_{1}=-\kappa_{2}=1.0$ takes place at the same value of the coupling strength $\varepsilon\simeq 1.0$. However, the entanglement of the ground state at the transition point is higher for the FP model. Notably, the FP model has permutation symmetry, but the nonzero torsional model with $\kappa_{1}=-\kappa_{2}=1.0$ does not have that symmetry. Therefore, the saturation values of the entanglement for these two models are different: the FP model shows spurious nonzero entanglement with $S_{V}\sim\ln 2$, and the other model becomes completely unentangled and then increases slowly with the coupling strength. The entanglement of the most excited state behaves qualitatively similarly for both these models. Now, for the other nonzero torsional case with $\kappa_{1}=\kappa_{2}=1.0$, the ground state entanglement makes a transition at a larger coupling strength $\varepsilon\simeq 2.0$ and then the entanglement reaches the saturation value $S_{V}=\ln 2$ around $\varepsilon\simeq 4.0$. The entanglement of the excited state of this model behaves similarly to its ground state entanglement, except for a sharp transition to the unentangled state around $\varepsilon\simeq 4.0$.
VII Quantum phase transition and Entanglement: Role of underlying classical dynamics
We now study the quantum phase transition (QPT) in the coupled top with torsion-free and nonzero torsion cases by observing transition in the ground state. Similarly, we study dynamical transition (DT) in these systems by observing transition in their excited state. We have related these transitions to the entanglement between the two tops and analyzed all the results from the underlying classical dynamics. Notably, we have shown how the system’s stable to unstable transition of the classical fixed point (CFP) influences the quantum transitions. We observe that the quantum transitions and the transition in the classical fixed points (CFPs) stability occur at the same coupling strength $\varepsilon$ for torsion-free and nonzero torsional models. The CFPs of the underlying classical dynamics of these systems describe their steady states, and these are determined by setting the time-derivatives of the dynamical variables $(Z_{i},\phi_{i})$ for $i=1,2$ at the left-hand side of Eq. (12) equal to zero. We now obtain a set of four coupled homogenous algebraic equations, and the solutions of these equations give the CFPs of the system for different coupling strengths $\varepsilon$. We find the same four CFPs for the torsion-free and nonzero torsional cases:
The generic behavior of the CFPs of the system is the following: Initially, for very weak coupling strengths $\varepsilon\gtrsim 0.0$, the system has two steady states (CFP-I and CFP-II), when both the spins of the coupled top are in the same direction ($X$-direction) with zero magnetization along $Z$-axis, as shown in Fig. 5. This figure schematically shows that the steady states CFP-I and CFP-II are distinguished from the direction of the spin of individual top: for CFP-I, both the spins are along negative $X$ direction, whereas for CFP-II case, the spins are along positive $X$-direction. These two steady states correspond to symmetry-unbroken stable fixed points. At some critical values of the coupling strength $\varepsilon=\varepsilon_{\rm c}$, the steady states become unstable and transition from the CFP-I state to the CFP-III state by bifurcation. However, the critical value $\varepsilon_{\rm c}$ is different for the steady states CFP-I and CFP-II. The CFP-I state bifurcates and forms a pair of symmetry-broken anti-ferromagnetic CFP-III states at the critical point. On the other hand, the CFP-II state bifurcates at a different critical point and forms a ferromagnetic steady state CFP-IV. These transitions are shown by a schematic diagram in Fig. 4.
We now present a detailed analysis of the QPT and DT for the torsion-free and nonzero torsional cases of the coupled top. In the left column of Fig. 4, we show the variation of the ground and the excited state energies with the coupling strength $\varepsilon$ to analyze the QPT and DT in the system. In the right column, we have shown the corresponding variation of the quantum entanglement between the two tops. However, here we focus on the QPT and its relation with the entanglement, so we have shown the entanglement for the coupling strengths $\varepsilon\in[0.0,3.0]$.
VII.1 FP model (torsion-free case)
In the FP model, Fig. 5(a) shows the ground and excited states energy variation with the coupling strength $\varepsilon$. Here, the bifurcation is observed in both the states. The ground state is stable at $E=-2.0$ when $\varepsilon<1.0$ (CFP-I state). Then, this state transitions to CFP-III state at $\varepsilon=1.0$, which remains stable for $\varepsilon>1.0$. In the CFP-III state, the fixed point varies with the coupling strength as the ground state energy varies with the coupling strength as $Z_{1}=-Z_{2}=\sqrt{1-\frac{1}{\varepsilon^{2}}}$ with $\phi_{1}=\phi_{2}=\pi$, and consequently, the ground state energy varies as $E=-(\varepsilon+\frac{1}{\varepsilon})$. This transition in the ground state describes the QPT in the system [39, 40].
On the other hand, for $\varepsilon<1.0$, the excited state energy of the system is stable at $E=+2$ (CFP-II state). This state transitions to the CFP-IV state by bifurcation at the critical point $\varepsilon=1.0$, and then it remains stable for $\varepsilon>1$. After the transition of the excited state from CFP-II to CFP-IV state, the fixed point also varies with $\varepsilon$ in the same fashion as the CFP-III state of the ground state. However, the excited state energy now increases as $E=(\varepsilon+\frac{1}{\varepsilon})$.
For the torsion-free case, we observe that the QPT and DT are happening simultaneously around $\varepsilon\simeq 1.0$. Figure 5(b) shows the variation in entanglement (von Neumann entropy) of the ground state (in red color) and excited state (in blue color) as a function of the coupling strength. We find that the entanglement of the ground state and the excited state maintain a steady growth till and then suddenly falls at $\varepsilon\simeq 1.0$. Interestingly, this transition in the entanglement is concurrent with QPT, DT, and the transition in the stability of the fixed points of the underlying classical dynamical system.
VII.2 Nonzero torsion case
We now present the results for two nonzero torsional cases: (a) both the torsion strengths are equal ($\kappa_{1}=\kappa_{2}=1.0$); and (b) they are equal in magnitude, but opposite in sign ($\kappa_{1}=-\kappa_{2}=1.0$).
VII.2.1 $\kappa_{1}=\kappa_{2}=1.0$
Figure 5(c) shows the variation in ground state and excited state energy as a function of the coupling strength $\varepsilon$. This figure shows that the CFP-I state becomes unstable at $\varepsilon=(\kappa_{1}+1)=2$ and then bifurcates into the CFP-III steady state. The energy of the CFP-III state varies with the coupling strength as $E=(1-\varepsilon)+\frac{1}{1-\varepsilon}$. On the other hand, the CFP-II state becomes unstable when $\varepsilon>(-\kappa_{1}+1)=0$. However, the effect of the unstable CFP-II state becomes prominent at a finite but very small value $\varepsilon\simeq 0.23$, where the CFP-II state bifurcates into the CFP-IV state. The energy of the CFP-IV state follows $E=(1+\varepsilon)+\frac{1}{(1+\varepsilon)}$. In Fig. 5(d), the transition in entanglement is again exactly happening at the point of bifurcation for the ground and excites states. However, unlike the previous case, the QPT and DT do not coincide here, but these transitions are consistent with the corresponding entanglement transition and the bifurcation of the steady states.
VII.2.2 $\kappa_{1}=-\kappa_{2}=1.0$
The results corresponding to this case are shown in Fig. 5(e) and (f). From Fig. 5(e), we observe that this system makes QPT and DT simultaneously at $\varepsilon\simeq 1.0$. This result is similar to the coupled top with the torsion-free case. Here, at the classical limit, the effect of torsion is canceled because of the opposite sign with the same magnitude of the torsion strength of each top. The energy for the CFP-III is calculated as $E=-\frac{2}{1-\varepsilon}-\varepsilon\left[1-\frac{1}{(1-\varepsilon)^{2}}\right]$; whereas for the CFP-IV state, the energy follows the relation $E=\frac{2}{1+\varepsilon}-\varepsilon\left[1-\frac{1}{(1+\varepsilon)^{2}}\right]$. Figure 5(f) shows the transition in entanglement, which is consistent with the QPT and DT in the system.
The above analysis reveals that the QPT and DT coincide around $\varepsilon\simeq 1.0$ in the coupled top with torsion-free (FP model) and nonzero torsion with $\kappa_{1}=-\kappa_{2}=1.0$. A common feature of these two models is that they both have chiral symmetry and thus follow one of the nonstandard symmetry classes. Since we only consider integer spin here, these models follow the BDI symmetry class. On the other hand, the coupled top with nonzero torsion and where $\kappa_{1}=-\kappa_{2}=1.0$ does not follow chiral symmetry. Interestingly, in this model, the QPT and DT do not coincide. However, these phase transitions are consistent with the corresponding entanglement transition and the transition in the stability of the fixed point of the underlying classical system.
VIII Conclusion
We have applied Floquet theory to a coupled kicked-top system at the high-frequency limit and have obtained an effective time-independent Hamiltonian employing a perturbation theory with the inverse of the frequency as the perturbation parameter. Then, the remaining part of the paper focuses on the classical and quantum properties of this time-independent Hamiltonian. In this system, when we neglect the nonlinear torsion of each top, the system becomes the well-known Feingold-Peres (FP) model. Besides the FP model, we also extensively study the coupled top model for two nonzero torsion (NZT) cases: torsion strengths of the individual top are equal (NZT-I), and torsion strengths are equal in magnitude but opposite in sign (NZT-II). The quantum version of the FP model and the NZT-II model follow the BDI symmetry class or chiral orthogonal symmetry class, which is one of the nonstandard symmetry classes. However, these two models with BDI symmetry are distinguishable from the permutation symmetry point of view: the first one (FP model) has this symmetry, but the other one (NZT-II) does not.
On the other hand, the NZT-I model has permutation symmetry but does not have chiral symmetry. Therefore, the NZT-I model does not follow any nonstandard symmetry class. Interestingly, we have obtained the NZT-II model from the NZT-I model by breaking the permutation symmetry via setting the torsion strengths of the two tops opposite in sign. The breaking of the permutation symmetry in the NZT-I model facilitates the construction of a chirality operator for the NZT-II model. We have also studied entanglement between the two tops as a function of the coupling strength by calculating the von Neumann entropy. The most important observation of this paper is that, for the systems with BDI symmetry class (FP and NZT-II), the QPT (the phase transition of the ground state), the DT (the phase transition of the ground state), and the transition in the entanglement happens simultaneously. However, in the NZT-I system, the QPT and the DT do not coincide; however, individually, these quantum transitions coincide with the corresponding entanglement transition. Interestingly, all these transitions are consistent with the stability of the fixed points of the underlying classical dynamical systems.

### v2

@@ Appendix A Equations of Motion
Following the prescription defined in Sec. III, we present a rigorous derivation of the equation of motion of only one component of the angular momentum, say $J_{x1}$. Others are derived identically, and we present the final form of them. Equation of motion of angular momentum operator ${J}_{x1}$ can be obtained as:
$\frac{d{J}_{x1}}{dt}$ $=-i[{J}_{x1},{H}_{\rm eff}]$
$=-\frac{i}{2j}\biggl{(}\kappa_{1}\,[{J}_{x1},{J}_{z1}^{2}]+2\varepsilon\,[{J}_{x1},{J}_{z1}]{J}_{z2}\biggr{)}$
$=-\frac{i}{2j}\biggl{\{}\kappa_{1}\,\bigl{(}[{J}_{x1},{J}_{z1}]{J}_{z1}+{J}_{z1}[{J}_{x1},{J}_{z1}]\bigr{)}$
@@
Now, divide the above equation by $j$ and obtain the form:
@@
Appendix B Phase space
In this section, we have plotted the phase space dynamics of NZT-I and NZT-II models for different coupling strength $\varepsilon$, around the two initial conditions shown in red and magenta circles in Fig. 2.
Figure 7: A bunch of trajectories is shown for the NZT-I model. (a)-(d): The trajectories are initiated around $Z_{1}=-0.7$ and $\Phi_{1}/\pi=0.1$ shown by a solid red circle in Fig. 2. (e)-(h): The trajectories are initiated around $Z_{1}=-0.1$ and $\Phi_{1}/\pi=0.1$, shown by a solid magenta circle in Fig. 2.
In Figs. 7(a)-(d), a bunch of trajectories are plotted for the initial conditions around $Z_{1}=-0.7$ and $\Phi_{1}/\pi=0.1$ for different coupling strengths. On the other hand, Figs. 7(e)-(h) show the trajectories initiated around another point $Z_{1}=-0.1$ and $\Phi_{1}/\pi=0.1$. Figures 7(a) and (e) suggest that, for the uncoupled case $\varepsilon=0$, identical regular islands are visible for both the initial conditions.
As the coupling strength increases $\varepsilon=0.3$, Fig. 7(b) indicates the beginning of transitions in the phase space. This figure shows that the observed single regular region for the uncoupled case is now bifurcating and forming two separate regular islands. This transition was detected in the Lyapunov exponent result presented in Fig. 3(a). However, we do not observe any bifurcation of the regular island for the other initial condition shown in Fig. 7(f). The corresponding Lyapunov exponent calculation in Fig. 3(b) also does not show any peaks, which is consistent with the behavior of the phase space trajectories.
We further increase the coupling strength $\varepsilon=1.0$. Figure 7(c) shows complete bifurcation of the initial single regular region into two regular regions. However, this figure shows one regular island at the left, because we have only considered the trajectories initiated around the red solid circle shown in Fig. 2. In 7(d), we do not see any further transition in the phase space trajectories for stronger coupling strength, except for a minute movement of the regular island. Therefore, the absence of any peaks in the Lyapunov exponent result in Fig. 3(a) for stronger coupling strengths is consistent with the behavior of the phase space trajectories. For the other initial conditions, the phase-space trajectories initiated from the region marked by the solid magenta circle in Fig. 2 show significantly different behavior. However, since the phase-space has the periodic property (i.e., left and right boundaries are the same line, whereas top and bottom boundaries are also the same line), a careful observation of Figs. 7(g) and (h) will reveal that each initial condition generates a closed loop on the surface of the sphere, and consequently, the trajectories are not sensitive to the initial conditions. This property is also reflected in the Lyapunov exponent result presented in Fig. 3(b).
Figure 8: A bunch of trajectories are shown for the NZT-II model. (a)-(d): The trajectories are initiated around $Z_{1}=-0.7$ and $\Phi_{1}/\pi=0.1$ shown by a solid red circle in Fig. 2. (e)-(h): The trajectories are initiated around $Z_{1}=-0.1$ and $\Phi_{1}/\pi=0.1$, shown by a solid magenta circle in Fig. 2.
We have also plotted many trajectories around the same two initial conditions for the NZT-II model in Fig. 8. For $\varepsilon=0$, regular islands are observed for both the initial conditions, as shown in Figs. 8(a) and (e). For $\varepsilon=0.5$, transitions in trajectories have been observed in both cases. In the NZT-II model, the bifurcation of the trajectories starts around $\varepsilon=0.5$. At $\varepsilon=1.0$, we see two prominent regular islands within a regular envelope trajectory observed in the uncoupled tops. However, unlike the single top, the trajectories start from one island and can reach another by the Arnold diffusion process, a typical phenomenon observed in higher-dimensional nonintegrable systems. For higher values of coupling strength $\varepsilon>1.5$, we do not see any qualitative change in the phase space trajectories. The corresponding Lyapunov exponent result shows the beginning of its steady growth at $\varepsilon=0.5$, reaching the maxima at $\varepsilon=1.0$, and then again goes down. Furthermore, we also conclude that the transition in phase space dynamics of NZT-I has been observed at $\varepsilon=0.3$, and for NZT-II it is observed at $\varepsilon=1.0$. These results are consistent with the observed transition in the entanglement between the two tops in the ground and most excited state energies.